OpenStudy (anonymous):

i need to find 2 distinct numbers where (x+3)/(x^2+9)=1/6

OpenStudy (ash2326):

We have\[\frac{x+3}{x^2+9}=\frac{1}{6}\] cross multiplying we get \[ 6x+18=x^2+9\] Bring all the terms to the same side \[x^2-6x-9=0\] Can you solve it now using quadratic formula?

OpenStudy (anonymous):

i believe so. 6+/-square root 36-4(1)(-9)/2

OpenStudy (ash2326):

yeah, correct :D

OpenStudy (anonymous):

huzzah

OpenStudy (anonymous):

confused though

OpenStudy (ash2326):

Why?

OpenStudy (anonymous):

what would the answer be? 6/-(36-36)/2?

OpenStudy (anonymous):

6+/-(0)/2

OpenStudy (ash2326):

\[x =\frac{6\pm\sqrt{36-4 \times (1)\times (-9) }}{2}\] we get \[x =\frac{6\pm\sqrt{72 }}{2}\] 72=36*2 so \[x =\frac{6\pm6\sqrt{2 }}{2}\] \[x=3\pm 3\sqrt{2}\] or \[x=3(1\pm \sqrt{2})\]

OpenStudy (anonymous):

my mistake, thanks ash